Patrick is an automobile salesman who receives a salary of $550.00 per week plus a commission of 5 percent on all sales. During a 4 week period he sold $45,500.00 worth of cars. What were Patricks average earnings?
Answer by Mathtut(3670)

45500/4=11375..avg sales per week
:
550+.05(11375)=550+568.75=$1018.75 avg earnings per week